How much was borrowed if a simple interest rate of 7.5% was offered payable in 12 months with interest of 105


1
Expert's answer
2022-01-31T16:17:20-0500

If a simple interest rate of 7.5% was offered payable in 12 months with interest of 105, then the sum which was borrowed is:

"PV\u00d7(1 + 0.075) = (PV + 105),"

PV = 105/0.075 = 1400.